OpenStudy (anonymous):

On Wednesday you invited 3 friends to a party. On Thursday each of these friends invited 3 friends. On Friday each one invited 3 friends. How many people were invited on Friday? Show your answer as a power. How many people were invited in all?

OpenStudy (anonymous):

Friday 3^3 (=27) In all 3^1+3^2+3^3 = 3+9+27=39